# Pixel Is Not A Barrier: An Effective Evasion Attack for Pixel-Domain Diffusion Models

### v1

@@ Appendix A More Implementation Details
$\gamma_{1}$ ($\mathcal{L}_{\text{attack}}$) $\gamma_{2}$ ($\mathcal{L}_{\text{fidelity}}$)
@@ B.1 Attack Effectiveness on Latent Diffusion Models
We propose the feature representation attacking loss which can be adapted to target any UNet-based diffusion models. Hence, it is applicable to attack LDM using our proposed framework. We follow the evaluation settings of the previous works (Xue et al. 2024b) for fair comparisons. Quantitative results are shown in Table 6. Compared to previous LDM-specified methods, our method could achieve comparable results. This finding reflects the general vulnerability in UNet-based diffusion models that can be exploited to craft adversarial images against either PDMs or LDMs.
@@
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\downarrow$ PSNR $\downarrow$ LPIPS $\uparrow$ IA-Score $\downarrow$
Church AdvDM ($+$) 0.85 $\pm$ 0.03 30.42 $\pm$ 0.15 0.23 $\pm$ 0.06 0.19 $\pm$ 0.05 28.00 $\pm$ 0.16 0.71 $\pm$ 0.04 0.49 $\pm$ 0.06
Mist ($+$) 0.81 $\pm$ 0.03 29.45 $\pm$ 0.13 0.25 $\pm$ 0.05 0.14 $\pm$ 0.03 27.95 $\pm$ 0.13 0.76 $\pm$ 0.04 0.48 $\pm$ 0.05
Diff-Protect (--) 0.79 $\pm$ 0.03 29.92 $\pm$ 0.15 0.24 $\pm$ 0.06 0.15 $\pm$ 0.03 28.00 $\pm$ 0.14 0.71 $\pm$ 0.04 0.48 $\pm$ 0.05
Diff-Protect ($+$) 0.79 $\pm$ 0.04 29.47 $\pm$ 0.11 0.26 $\pm$ 0.06 0.17 $\pm$ 0.04 28.00 $\pm$ 0.15 0.69 $\pm$ 0.04 0.49 $\pm$ 0.06
@@
Cat AdvDM ($+$) 0.86 $\pm$ 0.04 30.68 $\pm$ 0.24 0.25 $\pm$ 0.09 0.21 $\pm$ 0.05 28.03 $\pm$ 0.21 0.70 $\pm$ 0.07 0.53 $\pm$ 0.04
Mist ($+$) 0.81 $\pm$ 0.04 29.63 $\pm$ 0.22 0.27 $\pm$ 0.08 0.14 $\pm$ 0.04 27.96 $\pm$ 0.17 0.77 $\pm$ 0.06 0.52 $\pm$ 0.04
Diff-Protect (--) 0.78 $\pm$ 0.05 30.12 $\pm$ 0.24 0.27 $\pm$ 0.08 0.16 $\pm$ 0.05 27.96 $\pm$ 0.15 0.72 $\pm$ 0.06 0.52 $\pm$ 0.03
Diff-Protect ($+$) 0.77 $\pm$ 0.06 29.56 $\pm$ 0.16 0.28 $\pm$ 0.09 0.17 $\pm$ 0.05 27.98 $\pm$ 0.16 0.71 $\pm$ 0.06 0.53 $\pm$ 0.04
@@
Face AdvDM ($+$) 0.83 $\pm$ 0.02 30.81 $\pm$ 0.22 0.32 $\pm$ 0.06 0.26 $\pm$ 0.05 28.07 $\pm$ 0.28 0.74 $\pm$ 0.05 0.47 $\pm$ 0.07
Mist ($+$) 0.79 $\pm$ 0.03 29.75 $\pm$ 0.22 0.34 $\pm$ 0.06 0.19 $\pm$ 0.05 27.99 $\pm$ 0.21 0.81 $\pm$ 0.05 0.46 $\pm$ 0.08
Diff-Protect (--) 0.74 $\pm$ 0.04 30.34 $\pm$ 0.13 0.33 $\pm$ 0.06 0.21 $\pm$ 0.05 28.03 $\pm$ 0.21 0.76 $\pm$ 0.06 0.45 $\pm$ 0.07
Diff-Protect ($+$) 0.72 $\pm$ 0.05 29.68 $\pm$ 0.09 0.36 $\pm$ 0.06 0.21 $\pm$ 0.04 28.05 $\pm$ 0.22 0.74 $\pm$ 0.06 0.47 $\pm$ 0.07
@@
Table 6: Quantitative results in attacking LDM.
@@ B.3 Qualitative Results of Loss Ablation
B.4 Different Forward Time-step Sampling
@@
Figure 7: Qualitative results of optimizing different fixed diffusion forward steps.
B.5 More Qualitative Results
@@
B.6 Example of Loss Curves
Appendix C Limitations
Appendix D Societal Impacts
Appendix E Details of Our Proposed Algorithm
E.1 AtkPDM Algorithm without Latent Optimization
Algorithm 2 AtkPDM
1:  Input: Image to be protected $\mathbf{x}$, attack budget $\delta>0$, and step size $\gamma_{1},\gamma_{2}>0$
3:  while $L_{\text{attack}}$ not convergent do
4:     Sample timestep: $t\sim[0,T]$
5:     Sample noise: $\epsilon_{1},\epsilon_{2}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
6:     Compute original noisy sample: $\mathbf{x}_{t}\leftarrow\mathcal{F}(\mathbf{x},t,\epsilon_{1})$
7:     Compute adversarial noisy sample: $\mathbf{x}^{\operatornamewithlimits{adv}}_{t}\leftarrow\mathcal{F}(\mathbf{x}^{\operatornamewithlimits{adv}},t,\epsilon_{2})$
8:     Update $\mathbf{x}^{\operatornamewithlimits{adv}}$ by Gradient Descent: $\mathbf{x}^{\operatornamewithlimits{adv}}\leftarrow\mathbf{x}^{\operatornamewithlimits{adv}}-\gamma_{1}\operatornamewithlimits{sign}(\nabla_{\mathbf{x}^{\operatornamewithlimits{adv}}}\mathcal{L}_{\text{attack}}(\mathbf{x}^{\operatornamewithlimits{adv}}_{t},{\mathbf{x}_{t}}))$
10:        $\mathbf{x}^{\operatornamewithlimits{adv}}\leftarrow\mathbf{x}^{\operatornamewithlimits{adv}}-\gamma_{2}\nabla_{\mathbf{x}^{\operatornamewithlimits{adv}}}\mathcal{L}_{\text{fidelity}}(\mathbf{x}^{\operatornamewithlimits{adv}},\mathbf{x})$
11:     end while
12:  end while
13:  return  $\mathbf{x}^{\operatornamewithlimits{adv}}$
E.2 2-Wasserstein Distance Between Two Normal Distribution
@@ E.1 AtkPDM Algorithm without Latent Optimization
$\mathcal{W}_{2}^{2}(\mathcal{N}_{t},\mathcal{N}_{t}^{\operatornamewithlimits{adv}})=\min_{\pi\in\Pi(\mathcal{N}_{t},\mathcal{N}_{t}^{\operatornamewithlimits{adv}})}\int\|\mathbf{x}_{t}-\mathbf{x}_{t}^{\operatornamewithlimits{adv}}\|_{2}^{2}\text{d}\pi(\mathbf{x}_{t},\mathbf{x}_{t}^{\operatornamewithlimits{adv}}).$
@@
$\int\|\mu_{t}-\mu_{t}^{\operatornamewithlimits{adv}}\|_{2}^{2}\text{d}\pi(\mathbf{x}_{t},\mathbf{x}_{t}^{\operatornamewithlimits{adv}})=\|\mu_{t}-\mu_{t}^{\operatornamewithlimits{adv}}\|_{2}^{2},$
$\int\|\mathbf{x}_{t}-\mu_{t}\|_{2}^{2}\text{d}\pi(\mathbf{x}_{t},\mathbf{x}_{t}^{\operatornamewithlimits{adv}})=\operatornamewithlimits{trace}(\Sigma_{t}),$
$\int\|\mathbf{x}_{t}^{\operatornamewithlimits{adv}}-\mu_{t}^{\operatornamewithlimits{adv}}\|_{2}^{2}\text{d}\pi(\mathbf{x}_{t},\mathbf{x}_{t}^{\operatornamewithlimits{adv}})=\operatornamewithlimits{trace}(\Sigma_{t}^{\operatornamewithlimits{adv}}),$
$\int(\mathbf{x}_{t}-\mu_{t})^{\top}(\mathbf{x}_{t}^{\operatornamewithlimits{adv}}-\mu_{t}^{\operatornamewithlimits{adv}})\text{d}\pi(\mathbf{x}_{t},\mathbf{x}_{t}^{\operatornamewithlimits{adv}})$
$=\operatornamewithlimits{trace}\left(\mathbb{E}[(\mathbf{x}_{t}-\mu_{t})(\mathbf{x}_{t}^{\operatornamewithlimits{adv}}-\mu_{t}^{\operatornamewithlimits{adv}})^{\top}\right).$
@@
$C=\mathbb{E}\left[(\mathbf{x}_{t}-\mu_{t})(\mathbf{x}_{t}^{\operatornamewithlimits{adv}}-\mu_{t}^{\operatornamewithlimits{adv}})^{\top}\right].$
@@
$\text{maximum}\quad\operatornamewithlimits{trace}(C),$ (9)
@@ E.3 Alternating Optimization
$F(\mathbf{x},\mathbf{y})=F_{1}(\mathbf{x},\mathbf{y})+\lambda F_{2}(\mathbf{x},\mathbf{y}),$ (11)
where $\lambda>0$ is the Lagrange multiplier and $F_{1}$, $F_{2}$ are defined as
$F_{1}(\mathbf{x},\mathbf{y})$ $=\mathcal{L}_{\text{attack}}(\mathcal{F}(\mathbf{x},t,\epsilon_{1}),\mathcal{F}(\mathbf{y},t,\epsilon_{2})),$ (12)
$F_{2}(\mathbf{x},\mathbf{y})$ $=\max(\epsilon-\mathcal{L}_{\text{fidelity}}(\mathbf{x},\mathbf{y}),\mathbf{0}).$ (13)
@@
$\mathbf{y}^{i+\frac{1}{2}}=\operatornamewithlimits{argmin}_{\mathbf{y}}\left(F_{1}(\mathbf{x},\mathbf{y})+\lambda F_{2}(\mathbf{x},\mathbf{y}^{i})\right),$ (14)
$\mathbf{y}^{i+1}=\operatornamewithlimits{argmin}_{\mathbf{y}}\left(F_{1}(\mathbf{x},\mathbf{y}^{i+\frac{1}{2}})+\lambda F_{2}(\mathbf{x},\mathbf{y})\right).$ (15)
@@
$\mathbf{y}^{i+1/2}=\mathbf{y}^{i}-\gamma_{1}\operatornamewithlimits{sign}\left(\nabla_{\mathbf{y}^{i}}F_{1}(\mathbf{x},\mathbf{y}^{i})\right),$ (16)
@@
$\mathbf{y}^{i+1}=\mathbf{y}^{i+\frac{1}{2}}-\tilde{\gamma}_{2}\nabla_{\mathbf{y}^{i+\frac{1}{2}}}\lambda F_{2}(\mathbf{x},\mathbf{y}^{i+\frac{1}{2}})$ (17)
$=\mathbf{y}^{i+\frac{1}{2}}-\gamma_{2}\nabla_{\mathbf{y}^{i+\frac{1}{2}}}F_{2}(\mathbf{x},\mathbf{y}^{i+\frac{1}{2}}).$ (18)
Note that the gradient of $F_{2}$ can be derived as follows:
$\nabla_{\mathbf{y}}F_{2}(\mathbf{x},\mathbf{y})=\mathbb{I}_{\mathcal{C^{\prime}}}\odot\nabla_{\mathbf{x}^{\operatornamewithlimits{adv}}_{t}}\mathcal{L}_{\text{fidelity}}(\mathbf{x},\mathbf{y}),$
@@
Please note that after references, we also provide more results presented in Figures 6,  7,  8, and 9, please refer to subsequent pages.

### v3

@@ Appendix A More Implementation Details
$\gamma_{\text{attack}}$ $\gamma_{\text{fidelity}}$
@@ B.1 Attack Effectiveness on Latent Diffusion Models
We propose the feature representation attacking loss which can be adapted to target any UNet-based diffusion models. Hence, it is applicable to attack LDM using our proposed framework. We follow the evaluation settings of the previous work (Xue et al. 2023) for fair comparisons. Quantitative results are shown in Table 7. Compared to previous LDM-specified methods (Liang et al. 2023; Liang and Wu 2023; Xue et al. 2023), our method could achieve comparable results. This finding reflects the general vulnerability in UNet-based diffusion models that can be exploited to craft adversarial images against either PDMs or LDMs.
@@
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\downarrow$ PSNR $\downarrow$ LPIPS $\uparrow$ IA $\downarrow$
Church AdvDM 0.85 $\pm$ 0.03 30.42 $\pm$ 0.15 0.23 $\pm$ 0.06 0.19 $\pm$ 0.05 28.00 $\pm$ 0.16 0.71 $\pm$ 0.04 0.49 $\pm$ 0.06
Mist 0.81 $\pm$ 0.03 29.45 $\pm$ 0.13 0.25 $\pm$ 0.05 0.14 $\pm$ 0.03 27.95 $\pm$ 0.13 0.76 $\pm$ 0.04 0.48 $\pm$ 0.05
Diff-Protect 0.79 $\pm$ 0.03 29.92 $\pm$ 0.15 0.24 $\pm$ 0.06 0.15 $\pm$ 0.03 28.00 $\pm$ 0.14 0.71 $\pm$ 0.04 0.48 $\pm$ 0.05
@@
Cat AdvDM 0.86 $\pm$ 0.04 30.68 $\pm$ 0.24 0.25 $\pm$ 0.09 0.21 $\pm$ 0.05 28.03 $\pm$ 0.21 0.70 $\pm$ 0.07 0.53 $\pm$ 0.04
Mist 0.81 $\pm$ 0.04 29.63 $\pm$ 0.22 0.27 $\pm$ 0.08 0.14 $\pm$ 0.04 27.96 $\pm$ 0.17 0.77 $\pm$ 0.06 0.52 $\pm$ 0.04
Diff-Protect 0.78 $\pm$ 0.05 30.12 $\pm$ 0.24 0.27 $\pm$ 0.08 0.16 $\pm$ 0.05 27.96 $\pm$ 0.15 0.72 $\pm$ 0.06 0.52 $\pm$ 0.03
@@
Face AdvDM 0.83 $\pm$ 0.02 30.81 $\pm$ 0.22 0.32 $\pm$ 0.06 0.26 $\pm$ 0.05 28.07 $\pm$ 0.28 0.74 $\pm$ 0.05 0.47 $\pm$ 0.07
Mist 0.79 $\pm$ 0.03 29.75 $\pm$ 0.22 0.34 $\pm$ 0.06 0.19 $\pm$ 0.05 27.99 $\pm$ 0.21 0.81 $\pm$ 0.05 0.46 $\pm$ 0.08
Diff-Protect 0.74 $\pm$ 0.04 30.34 $\pm$ 0.13 0.33 $\pm$ 0.06 0.21 $\pm$ 0.05 28.03 $\pm$ 0.21 0.76 $\pm$ 0.06 0.45 $\pm$ 0.07
@@
Table 7: Quantitative results in attacking LDM. The best is marked in bold and the second best is underlined. Errors denote one standard deviation of all images in our test datasets.
@@ B.3 Qualitative Results of Loss Ablation
B.4 Example of Loss Curves
B.5 Different Forward Time-step Sampling
@@
Figure 8: Qualitative results of optimizing different fixed diffusion forward steps $t^{*}$.
B.6 More Qualitative Results
@@
Appendix C Backgrounds of Diffusion Models
Score-based models and diffusion models allowing generate samples starting from easy-to-sample Gaussian noise to complex target distributions. Starting from Gaussian noise, the sampling process iteratively applies the score function, i.e., $\nabla_{\mathbf{x}}\log p(\mathbf{x})$ of the complex target distribution $p(\mathbf{x})$ to generate the sample from $p(\mathbf{x})$. The exact estimation of the ground truth score function is intractable since the score function is the derivative of the target distribution $p(\mathbf{x})$. However, we can approximate the score function without directly dealing with $p(\mathbf{x})$. Song et al. proposed score-based models (Song et al. 2020) to learn the score function effectively via score matching. Ho et al. proposed Denoising Diffusion Probability Model (DDPM) (Ho, Jain, and Abbeel 2020), providing another perspective on learning score function with noise perturbed data, allowing more effective low-density area estimation and improving the mode diversity, thereby capable of generating highly sophisticated data, e.g., natural images. In a nutshell, training DDPM involves perturbing data with Gaussian noise in different timestep-controlled variance schedules, i.e., forward diffusion, and a parametrized model $\epsilon_{\theta}(\mathbf{x}_{t},t)$ will learn to predict the added noise conditioning on noisy data $\mathbf{x}_{t}$ and current noise level $t$. Sampling with learned DDPM starts with random noise and iteratively applies the model $\epsilon_{\theta}(\mathbf{x}_{t},t)$ to denoise, i.e., reverse diffusion sampling, thereby generating a sample from the learned distribution. Specifically, for forward diffusion, we perturb the data with a linear combination of Gaussian noise and clean data as $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$ via a scheduler $\bar{\alpha}_{t}$ controlling the strength of added noise, here $t\in[0,T]$ and $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, note that when $t$ reaches $T$, the perturbed data $\mathbf{x}_{t}$ become Gaussian noise. The training objective of the $\epsilon_{\theta}$ is defined as the noise prediction MSE $\mathbb{E}_{t,\mathbf{x},\epsilon_{t}}[\|\epsilon_{t}-\epsilon_{\theta}(\mathbf{x}_{t},t)\|_{2}^{2}]$. In sampling with diffusion models, Song et.al proposed DDIM (Song, Meng, and Ermon 2021) that generalized the DDPM sampling formulation as:
$\mathbf{x}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)$
The first term of the right-hand side of the equation represents direct clean sample estimation $\hat{\mathbf{x}}_{0}$ from noisy sample $\mathbf{x}_{t}$ which is derived from Tweedie’s formula(Efron 2011). Therefore, the noise prediction can bridge with the score function via Tweedie’s formula, where the denoising objective and score-matching objective are identical.
Appendix D Details of Our Proposed Algorithm
D.1 2-Wasserstein Distance Between Two Normal Distribution
@@ Appendix D Details of Our Proposed Algorithm
$\mathcal{W}_{2}^{2}(\mathcal{N}_{t},\mathcal{N}_{t}^{\operatornamewithlimits{adv}})=\min_{\pi\in\Pi(\mathcal{N}_{t},\mathcal{N}_{t}^{\operatornamewithlimits{adv}})}\int\|\mathbf{f}_{t}-\mathbf{f}_{t}^{\operatornamewithlimits{adv}}\|_{2}^{2}\text{d}\pi(\mathbf{f}_{t},\mathbf{f}_{t}^{\operatornamewithlimits{adv}}).$
@@
$\int\|\mu_{t}-\mu_{t}^{\operatornamewithlimits{adv}}\|_{2}^{2}\text{d}\pi(\mathbf{f}_{t},\mathbf{f}_{t}^{\operatornamewithlimits{adv}})=\|\mu_{t}-\mu_{t}^{\operatornamewithlimits{adv}}\|_{2}^{2},$
$\int\|\mathbf{f}_{t}-\mu_{t}\|_{2}^{2}\text{d}\pi(\mathbf{f}_{t},\mathbf{f}_{t}^{\operatornamewithlimits{adv}})=\operatornamewithlimits{trace}(\Sigma_{t}),$
$\int\|\mathbf{f}_{t}^{\operatornamewithlimits{adv}}-\mu_{t}^{\operatornamewithlimits{adv}}\|_{2}^{2}\text{d}\pi(\mathbf{f}_{t},\mathbf{f}_{t}^{\operatornamewithlimits{adv}})=\operatornamewithlimits{trace}(\Sigma_{t}^{\operatornamewithlimits{adv}}),$
$\int(\mathbf{f}_{t}-\mu_{t})^{\top}(\mathbf{f}_{t}^{\operatornamewithlimits{adv}}-\mu_{t}^{\operatornamewithlimits{adv}})\text{d}\pi(\mathbf{f}_{t},\mathbf{f}_{t}^{\operatornamewithlimits{adv}})$
$=\operatornamewithlimits{trace}\left(\mathbb{E}[(\mathbf{f}_{t}-\mu_{t})(\mathbf{f}_{t}^{\operatornamewithlimits{adv}}-\mu_{t}^{\operatornamewithlimits{adv}})^{\top}\right).$
@@
$C=\mathbb{E}\left[(\mathbf{f}_{t}-\mu_{t})(\mathbf{f}_{t}^{\operatornamewithlimits{adv}}-\mu_{t}^{\operatornamewithlimits{adv}})^{\top}\right].$
@@
$\text{maximum}\quad\operatornamewithlimits{trace}(C),$
@@ D.2 Alternating Optimization
$F(\mathbf{x},\mathbf{y})=F_{\text{attack}}(\mathbf{x},\mathbf{y})+\lambda F_{\text{fidelity}}(\mathbf{x},\mathbf{y}),$
where $\lambda>0$ is the Lagrange multiplier and $F_{\text{attack}}$, $F_{\text{fidelity}}$ are defined as
$F_{\text{attack}}(\mathbf{x},\mathbf{y})$ $=-\mathcal{L}_{\text{attack}}(\mathcal{F}(\mathbf{x},t,\epsilon),\mathcal{F}(\mathbf{y},t,\epsilon^{\operatornamewithlimits{adv}})),$
$F_{\text{fidelity}}(\mathbf{x},\mathbf{y})$ $=\max(\epsilon-\mathcal{L}_{\text{fidelity}}(\mathbf{x},\mathbf{y}),\mathbf{0}).$
@@
$\mathbf{y}^{i+\frac{1}{2}}=\operatornamewithlimits{argmin}_{\mathbf{y}}\left(F_{\text{attack}}(\mathbf{x},\mathbf{y})+\lambda F_{\text{fidelity}}(\mathbf{x},\mathbf{y}^{i})\right),$ (7)
$\mathbf{y}^{i+1}=\operatornamewithlimits{argmin}_{\mathbf{y}}\left(F_{\text{attack}}(\mathbf{x},\mathbf{y}^{i+\frac{1}{2}})+\lambda F_{\text{fidelity}}(\mathbf{x},\mathbf{y})\right).$ (8)
@@
$\mathbf{y}^{i+1/2}=\mathbf{y}^{i}-\gamma_{\text{attack}}\operatornamewithlimits{sign}\left(\nabla_{\mathbf{y}^{i}}F_{\text{attack}}(\mathbf{x},\mathbf{y}^{i})\right).$
@@
$\mathbf{y}^{i+1}=\mathbf{y}^{i+\frac{1}{2}}-\tilde{\gamma}_{\text{fidelity}}\nabla_{\mathbf{y}^{i+\frac{1}{2}}}\lambda F_{\text{fidelity}}(\mathbf{x},\mathbf{y}^{i+\frac{1}{2}})$
$=\mathbf{y}^{i+\frac{1}{2}}-\gamma_{\text{fidelity}}\nabla_{\mathbf{y}^{i+\frac{1}{2}}}F_{\text{fidelity}}(\mathbf{x},\mathbf{y}^{i+\frac{1}{2}}).$
Note that the gradient of $F_{\text{fidelity}}$ can be derived as follows:
$\nabla_{\mathbf{y}}F_{\text{fidelity}}(\mathbf{x},\mathbf{y})=\mathbb{I}_{\mathcal{C^{\prime}}}\cdot\nabla_{\mathbf{x}^{\operatornamewithlimits{adv}}_{t}}\mathcal{L}_{\text{fidelity}}(\mathbf{x},\mathbf{y}),$
@@
Please note that after references, we also provide more results presented in Figures 7,  8,  9, and 5, please refer to subsequent pages.
D.3 AtkPDM Algorithm without Latent Optimization
Algorithm 2 AtkPDM
1:  Input: Image to be protected $\mathbf{x}$, attack budget $\delta>0$, and step size $\gamma_{\text{attack}},\gamma_{\text{fidelity}}>0$
3:  while $L_{\text{attack}}$ not convergent do
4:     Sample timestep: $t\sim[0,T]$
5:     Sample noise: $\epsilon,\epsilon^{\operatornamewithlimits{adv}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
6:     Compute original noisy sample: $\mathbf{x}_{t}\leftarrow\mathcal{F}(\mathbf{x},t,\epsilon)$
7:     Compute adversarial noisy sample: $\mathbf{x}^{\operatornamewithlimits{adv}}_{t}\leftarrow\mathcal{F}(\mathbf{x}^{\operatornamewithlimits{adv}},t,\epsilon^{\operatornamewithlimits{adv}})$
8:     Update $\mathbf{x}^{\operatornamewithlimits{adv}}$ by Gradient Descent: $\mathbf{x}^{\operatornamewithlimits{adv}}\leftarrow\mathbf{x}^{\operatornamewithlimits{adv}}-\gamma_{\text{attack}}\operatornamewithlimits{sign}(\nabla_{\mathbf{x}^{\operatornamewithlimits{adv}}}(-\mathcal{L}_{\text{attack}}(\mathbf{x}^{\operatornamewithlimits{adv}}_{t},{\mathbf{x}_{t}})))$
10:        $\mathbf{x}^{\operatornamewithlimits{adv}}\leftarrow\mathbf{x}^{\operatornamewithlimits{adv}}-\gamma_{\text{fidelity}}\nabla_{\mathbf{x}^{\operatornamewithlimits{adv}}}\mathcal{L}_{\text{fidelity}}(\mathbf{x}^{\operatornamewithlimits{adv}},\mathbf{x})$
11:     end while
12:  end while
13:  return  $\mathbf{x}^{\operatornamewithlimits{adv}}$
Appendix E Limitations
Appendix F Societal Impacts